Find the formula for the focal length of a thin lens if the R2 radius of curvature

of the lens surface is k times more than R1 radius of curvature and the refraction index of

the lens is n = 2. The medium surrounding the lens is air.


1
Expert's answer
2022-01-11T09:04:33-0500
"\dfrac 1f=(n-1)\bigg(\dfrac 1{R_1}-\dfrac 1{R_2}\bigg),\\\space\\
\dfrac 1f=(2-1)\bigg(\dfrac 1{R_1}-\dfrac 1{kR_1}\bigg)=\dfrac{k-1}{kR_1},\\\space\\
f=\dfrac{kR_1}{k-1}."